OpenStudy (anonymous):

2. Is 4x-3= 19 true, false, or open? a. true. b. false c. open

OpenStudy (anonymous):

Is 4x-3= 19 true, false, or open? 4x-3= 19 4x-3 + 3 = 19 + 3 4x = 22 4x/4 = 22/4 x = 22/4 Open

OpenStudy (anonymous):

okay thank you :)